OpenStudy (anonymous):

***PLEASE HELP*** A cylinder has a radius of 1 cm and a height of 12 cm. The surface area of the cylinder is _____ cm2.

OpenStudy (anonymous):

37.71cm^2 (37.71 cm squared)

OpenStudy (anonymous):

actually 37.69cm2 pi × r 2 × h = 3.14159265358979 x1^2x12= 37.6991118430775

OpenStudy (anonymous):

my answer choices are - 26Pi 24Pi 30Pi 28Pi

OpenStudy (anonymous):

its 26Pi my apologies I calculated the area of a cylinder originally, the formula is 2xPi x r2 + 2xPi x r x h, = 2 x 3.14 x 1 x 1 + 2 x 3.14 x 1 x 12 = 6.28318 + 75.3982236 = 81.6814036 = 81.6814036 / Pi = 26Pi check your answer, Pi x 26 = 81.68140 xxx